OpenStudy (anonymous):

I'm stuck...... Solve the following systems of equations using matrices. x+y-5z=-18 3x-3y+z=6 x+3y-2z=-13 step by step please -_-

OpenStudy (anonymous):

calculator or reduced ?

OpenStudy (anonymous):

the answers in ordered pair form. (x,y,z)

OpenStudy (anonymous):

are you sure all your numbers and signs are correct? According to http://www.wolframalpha.com/input/?i=x%2By-5z%3D-18+3x-3y%2Bz%3D6+x%2B3y-2z%3D-13 all answers are fractions.

OpenStudy (agreene):

\[\left[\begin{matrix}1 & 1 & -5 \\ 3 & -3&1 \\ 1 & 3 &-2\end{matrix}\right]= \left[\begin{matrix} -18\\ 6 \\ -13 \end{matrix}\right]\]

OpenStudy (anonymous):

hmm how to solve how to solve.... agreene do kramers !!

OpenStudy (agreene):

lol, f that!

OpenStudy (anonymous):

XD kramers is much easier than say...... doing reduced maybe?

OpenStudy (agreene):

I would never solve any system by matrix unless I were forced to do so.

OpenStudy (anonymous):

since they're fractions? lol

OpenStudy (anonymous):

throwing them into the calculator is the easier XD

OpenStudy (anonymous):

where are you stuck nancy?

OpenStudy (agreene):

x+y-5z=-18 3x-3y+z=6 x+3y-2z=-13 ------------------------------ x=5z-18-y ---- 5z-18-y+3y-2z=-13 y= 1/2 ( 5-3z) ------ 3(5z-18-y)-3( 1/2 ( 5-3z)) +z = 6 3(5z-18-(1/2 ( 5-3z)))-3( 1/2 ( 5-3z)) +z = 6 z=3 ----- y= 1/2 ( 5-3z) y= 1/2 ( 5-3(3)) y=-2 ------ x=5z-18-y x=5(3)-18-(-2) x=-1 Solution: (-1,-2,3) Matrices, HA!

OpenStudy (agreene):

@MarinaDL, please learn how logic works: http://www.wolframalpha.com/input/?i=x%2By-5z%3D-18%3B+3x-3y%2Bz%3D6%3B+x%2B3y-2z%3D-13

OpenStudy (anonymous):

ouch cold hearted bro, he's out for blood

OpenStudy (agreene):

^_^

OpenStudy (agreene):

I'm a touch drunk atm.

OpenStudy (anonymous):

drunken mathematics = awesome

OpenStudy (agreene):

How do you think I survived getting a BA in Math? By being sober? LOL!

OpenStudy (anonymous):

drinking while helping with math probably makes you act like a raging genius. "How did you not know that this limit goes to zero. are you that dull "

OpenStudy (anonymous):

lol a drunk mathematician could be a comedy act

OpenStudy (agreene):

lol probably!

OpenStudy (anonymous):

I think we have something here

OpenStudy (agreene):

But yeah, I would have not survived my Topics classes without Rum/Gin/Beer... it was all non-euclidan geometry/topology... which is REALLY interesting, but also so mind numbingly hard that it makes you question your own existence as an organism with self thought.

OpenStudy (agreene):

spelling fails all up and in that post....

OpenStudy (anonymous):

whats the derivative of 2xcos(pi x) too lazyyy lol

OpenStudy (agreene):

\[\frac{d}{dx} ( 2x\cos(\pi x))=2 (\cos(π x)-π x \sin(π x))\] Product rule - > chain rule kid!

OpenStudy (anonymous):

i meant integral lmfao sooo tired

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

\[\int\limits_{-\pi}^{\pi}x^2+2xcos \pi x+\cos^2 \pi x\]

OpenStudy (agreene):

Why do I deal with you?

OpenStudy (agreene):

@outkast is that really what you mean to integrate? because that can be reduced... by a lot

OpenStudy (anonymous):

To agreene: thanks for the example. I figured out it wasn't right, but didn't know how to put it proper way.

OpenStudy (agreene):

Wait no, I was looking at it wrong, you mean that pi and x are in the cos function... eff that noise! @MarinaDL, No problem <3

OpenStudy (anonymous):

yesh i have to integrate this lol

OpenStudy (agreene):

you realize how annoying that integral will be right?

OpenStudy (agreene):

Like... you need to factor it, and then I see atleast 1 integration by parts followed by a trig sub... who knows after that, lol...

OpenStudy (agreene):

haha! Wolfram knows, and shows steps :D http://www.wolframalpha.com/input/?i=x%5E2%2B2xcos+%28%5Cpi+x%29%2B%5Ccos%5E2+%28%5Cpi+x%29+dx

OpenStudy (anonymous):

trig sub?... haha that isn't even all the question

OpenStudy (agreene):

wolfram used trig sub ^_^ It's super effective :D **not the sqrt(1-x^2)=asin(theta) trig sub... but the let u=random trig stuffs trig sub.

OpenStudy (anonymous):

@ ohsnapsitsnancy A solution to the equations is attached.